Let u = 2i – 3j + 4k v = 3i + k – 2k w = i+ 5j+3k

a. 2u + 3v + 4w

b. u . v and u . w

c. u x v and || u x v ||

d. Find the unit vector of u, v and w

e. Proj (u,w)

f. Establish the angle between the u and v.


1
Expert's answer
2021-11-08T16:57:32-0500

(a) "2u+3v+4w = 2(2i \u2013 3j + 4k) + 3(3i + j \u2013 2k) + 4(i + 5j + 3k) = 4i-6j+8k+9i+3j-6k+4i+20j+12k = 17i+17j+14k"


(b) "u\u2022v = (2i \u2013 3j + 4k)\u2022(3i + j \u2013 2k) = 6-3-8=-5"


"u\u2022w = (2i \u2013 3j + 4k)\u2022(i + 5j + 3k) = 2-15+12=-1"



(c) "u×v = \begin{pmatrix}
i & j & k \\
2 & -3 & 4 \\
3 & 1 & -2
\end{pmatrix} = i(6-4)-j(-4-12)+k(2+9) = 2i+16j+11k"


"|u×v| = \sqrt{2²+16²+11²} = \sqrt{381}"



(d) unit vector of u"= \frac{u}{|u|} = \frac{2i – 3j + 4k}{\sqrt{2²+(-3)²+4²}} = \frac{2i – 3j + 4k}{\sqrt{29}}"


unit vector of v = "\frac{v}{|v|} = \frac{3i + j – 2k}{\sqrt{3²+1²+(-2)²}} = \frac{3i + j – 2k}{\sqrt{14}}"


unit vector of w = "\frac{w}{|w|} = \frac{i + 5j + 3k}{\sqrt{1²+5²+3²}} = \frac{i + 5j + 3k}{\sqrt{35}}"


(e) Proj(u,w) = "\frac{u•w}{|w|²} w = \frac{-1}{35} (i + 5j + 3k)"


(f) Angle between u and v

"Cos∅ = \frac{u•v}{|u||v|} = \frac{-5}{\sqrt{29×14}} = \frac{-5}{\sqrt{406}}"


"=> ∅ = cos^{-1}(\frac{-5}{\sqrt{406}})"


"\u2205 = 104.37\u00b0"